OpenStudy (anonymous):

How do I find the exact value for cot(-arcsin(1/7))?

OpenStudy (anonymous):

you know what cot(-u) equals to?

OpenStudy (anonymous):

would it not be the same as -cot(u)?

OpenStudy (anonymous):

that's correct. so you have -cot(arcsin(1/7)) yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if you let θ = arcsin(1/7), what is the equivalent form?

OpenStudy (anonymous):

7?

OpenStudy (anonymous):

try again

OpenStudy (anonymous):

sin(theta) = 1/7?

OpenStudy (anonymous):

Im just not sure. I know cot(arcsin(x)) = sqrt{1-x^2/x}

OpenStudy (anonymous):

well, you can use that formula or you can through the process by hand. Either way will give u the correct answer

OpenStudy (anonymous):

it's supposed to be (sqrt(1-x^2))/x btw

OpenStudy (anonymous):

When I use the formula I get to (sqrt(48/49))/-1/7, but get stuck right there

OpenStudy (anonymous):

sqrt(48/49) = (sqrt(48)/7) / (-1/7) = -sqrt(48) = -4sqrt(3)

OpenStudy (anonymous):

Which would be the exact value?

OpenStudy (anonymous):

How would I find it by hand?

OpenStudy (anonymous):

-4sqrt(3)

OpenStudy (anonymous):

Yes, what is the method by hand?

OpenStudy (anonymous):

sin(θ) = 1/7  what's cot(θ) then?

OpenStudy (anonymous):

x/y which would give me the same answer. Where I could find x by using the Pythagorean. Awesome, much more simple than I thought. Thank you

OpenStudy (anonymous):

yw